On January 1, 2019, Romero Company purchased equipment for $265,000.
The equipment was assigned a $7,000 residual value and a 16-year
life. Romero Company will use the straight-line method to depreciate
the equipment. On January 1, 2028, Romero Company spent $29,725 to
overhaul the equipment. This capital expenditure resulted in Romero
Company changing the life of the equipment from 16 years to 22 years
and adjusting the residual value to be $4,000 at the end of the 22
years.

Calculate the depreciation expense recorded on the equipment for 2028.
Answer #1
Annual Depreciation on Equipment 16125 =(265000-7000)/16
Years of Equipment used 9 2019 to 2027
Accumulated Depreciation on Equipment 145125 =16125*9
Book value of Equipment on Jan 1,2028 119875 =265000-145125
Cost for Depreciation 145600 =119875+29725-4000
Divide by Remaining useful life 13 =22-9
Depreciation expense on the equipment for 2028 11200